If Don Gecewicz invests ​$3000 at 8​% interest compounded​ quarterly, find the amount after 4 years.
Answer by Boreal(15235)

Compound interest formula used here
P=Po(1+r/n)^nt, r is rate, n is number of compoundings in a year, t is the number of years
P=3000(1+0.08/4)^16
P=3000*1.02^16=$4118.36 ANSWER